123qwe:

Calculate the frequency of the n = 5 line in the Lyman series of hydrogen.

sillybilly123:

from Rydberg formula \( \dfrac{1}{ \lambda }=R \left(1-{\dfrac {1}{n^{2}}}\right) \), then: \( f = c R \left(1-{\dfrac {1}{n^{2}}}\right) \) \( f_{\delta} \approx 3 \cdot 10^ 8 \times 1.0968 \cdot10^{7}\left(1-{\dfrac {1}{25}}\right) = 316 \text{ THz}\) It's very sensitive to rounding....

Vocaloid:

Moved to Chem